You put a bottle of softdrink in a refrigerator and leave it until its temperature has dropped 10.0K.What is its temperature change in(a)Foand(b)Co?



1
Expert's answer
2021-06-24T06:03:30-0400

oC has the same size as the Kelvin, hence the temperature change is 10.0oC

"\degree{F}=\frac{9}{5}\times\degree{C}+32=\frac{9}{5}\times10.0+32=50\degree{F}"